OpenStudy (anonymous):

Solve the equation. 2x – 10 = 8

OpenStudy (anonymous):

again the same thing put the numbers on 1 side and the x on the other add 10 to both sides 2x-10+10=8+10 2x=18 divide by 2 x=9